Predict the products of the following reactions
1 equivalent HBr NaNH2 Br (excess) 1. Sia BH, THE 2. H2O2, NaOH HgSO4 H2SO4 H20
Answer #1

CC-H Br 1 equivalent c=CH₂ Hor Mechanism CECH Cc=c-н H- Br - CH₂ Br (=CH₂ Br g. Br Na NH₂ excess Erc=c-н Mechanism: CH NH₂ -sia, BH, THE H-CEC- HCCH H₂O₃, NaOH Mechanism: - се с Sa Sia - B -H B - C - THE UI sia sia H₂O2 H-O- NaoH Sia s o -c= BosiaHgSO4 H₂SO4, H₂ H₃c-en CH₂ CH2 CH₂ Mechanism: Н - С=С 7 7 H-c=0 Ha a mercuric enol | 4300 4 Hein